OpenStudy (iwanttogotostanford):

Solve 4^(2x) = 7^(x−1). 2.35389 −2.35389 −1 1

OpenStudy (iwanttogotostanford):

@peachpi

OpenStudy (anonymous):

do you have to do this by hand?

OpenStudy (iwanttogotostanford):

nope

OpenStudy (iwanttogotostanford):

i can use a calculator and stuff

OpenStudy (anonymous):

you'd have to take a log of both sides. then rewrite ln (4^(2x)) = ln (7^(x-1)) 2x ln 4 = (x - 1) ln 7

OpenStudy (anonymous):

oh if you can use a calculator, use wolfram alpha. or graph both sides and find the intersection point

OpenStudy (iwanttogotostanford):

ok thanks!!! I have one more and I am really confudes... I will medal you and everything :-) thanks so much for the help so far